# IP-LLM: Semantic Space Informed Prompt Learning with LLM for Time Series Forecasting

### v1

@@ Appendix D Full Few-shot Forecasting Results
| Methods | IP-LLM | Time-LLM | OFA | iTransformer | Dlinear | PatchTST | TimesNet | FEDformer | Autoformer | Stationary | ETSformer | LightTS | |||||||||||||
| MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | ||
| Weather | 96 | 0.154 | 0.201 | 0.161 | 0.210 | 0.163 | 0.215 | 0.253 | 0.307 | 0.171 | 0.224 | 0.165 | 0.215 | 0.184 | 0.230 | 0.188 | 0.253 | 0.221 | 0.297 | 0.192 | 0.234 | 0.199 | 0.272 | 0.217 | 0.269 |
| 192 | 0.195 | 0.241 | 0.204 | 0.248 | 0.210 | 0.254 | 0.292 | 0.328 | 0.215 | 0.263 | 0.210 | 0.257 | 0.245 | 0.283 | 0.250 | 0.304 | 0.270 | 0.322 | 0.269 | 0.295 | 0.279 | 0.332 | 0.259 | 0.304 | |
| 336 | 0.260 | 0.302 | 0.261 | 0.302 | 0.256 | 0.292 | 0.322 | 0.346 | 0.258 | 0.299 | 0.259 | 0.297 | 0.305 | 0.321 | 0.312 | 0.346 | 0.320 | 0.351 | 0.370 | 0.357 | 0.356 | 0.386 | 0.303 | 0.334 | |
| 720 | 0.303 | 0.329 | 0.309 | 0.332 | 0.321 | 0.339 | 0.365 | 0.374 | 0.320 | 0.346 | 0.332 | 0.346 | 0.381 | 0.371 | 0.387 | 0.393 | 0.390 | 0.396 | 0.441 | 0.405 | 0.437 | 0.448 | 0.377 | 0.382 | |
| Avg | 0.228 | 0.268 | 0.234 | 0.273 | 0.238 | 0.275 | 0.308 | 0.338 | 0.241 | 0.283 | 0.242 | 0.279 | 0.279 | 0.301 | 0.284 | 0.324 | 0.300 | 0.342 | 0.318 | 0.323 | 0.318 | 0.360 | 0.289 | 0.322 | |
| Electricity | 96 | 0.142 | 0.243 | 0.139 | 0.241 | 0.139 | 0.237 | 0.154 | 0.257 | 0.150 | 0.253 | 0.140 | 0.238 | 0.299 | 0.373 | 0.231 | 0.323 | 0.261 | 0.348 | 0.420 | 0.466 | 0.599 | 0.587 | 0.350 | 0.425 |
| 192 | 0.163 | 0.260 | 0.151 | 0.248 | 0.156 | 0.252 | 0.171 | 0.272 | 0.164 | 0.264 | 0.160 | 0.255 | 0.305 | 0.379 | 0.261 | 0.356 | 0.338 | 0.406 | 0.411 | 0.459 | 0.620 | 0.598 | 0.376 | 0.448 | |
| 336 | 0.173 | 0.270 | 0.169 | 0.270 | 0.175 | 0.270 | 0.196 | 0.295 | 0.181 | 0.282 | 0.180 | 0.276 | 0.319 | 0.391 | 0.360 | 0.445 | 0.410 | 0.474 | 0.434 | 0.473 | 0.662 | 0.619 | 0.428 | 0.485 | |
| 720 | 0.237 | 0.320 | 0.240 | 0.322 | 0.233 | 0.317 | 0.263 | 0.348 | 0.223 | 0.321 | 0.241 | 0.323 | 0.369 | 0.426 | 0.530 | 0.585 | 0.715 | 0.685 | 0.510 | 0.521 | 0.757 | 0.664 | 0.611 | 0.597 | |
| Avg | 0.178 | 0.273 | 0.175 | 0.270 | 0.176 | 0.269 | 0.196 | 0.293 | 0.180 | 0.280 | 0.180 | 0.273 | 0.323 | 0.392 | 0.346 | 0.427 | 0.431 | 0.478 | 0.444 | 0.480 | 0.660 | 0.617 | 0.441 | 0.489 | |
| Traffic | 96 | 0.401 | 0.285 | 0.418 | 0.291 | 0.414 | 0.297 | 0.448 | 0.329 | 0.419 | 0.298 | 0.403 | 0.289 | 0.719 | 0.416 | 0.639 | 0.400 | 0.672 | 0.405 | 1.412 | 0.802 | 1.643 | 0.855 | 1.157 | 0.636 |
| 192 | 0.410 | 0.293 | 0.414 | 0.296 | 0.426 | 0.301 | 0.487 | 0.360 | 0.434 | 0.305 | 0.415 | 0.296 | 0.748 | 0.428 | 0.637 | 0.416 | 0.727 | 0.424 | 1.419 | 0.806 | 1.641 | 0.854 | 1.207 | 0.661 | |
| 336 | 0.425 | 0.314 | 0.421 | 0.311 | 0.434 | 0.303 | 0.514 | 0.372 | 0.449 | 0.313 | 0.426 | 0.304 | 0.853 | 0.471 | 0.655 | 0.427 | 0.749 | 0.454 | 1.443 | 0.815 | 1.711 | 0.878 | 1.334 | 0.713 | |
| 720 | 0.470 | 0.330 | 0.462 | 0.327 | 0.487 | 0.337 | 0.532 | 0.383 | 0.484 | 0.336 | 0.474 | 0.331 | 1.485 | 0.825 | 0.722 | 0.456 | 0.847 | 0.499 | 1.539 | 0.837 | 2.660 | 1.157 | 1.292 | 0.726 | |
| Avg | 0.426 | 0.305 | 0.429 | 0.306 | 0.440 | 0.310 | 0.495 | 0.361 | 0.447 | 0.313 | 0.430 | 0.305 | 0.951 | 0.535 | 0.663 | 0.425 | 0.749 | 0.446 | 1.453 | 0.815 | 1.914 | 0.936 | 1.248 | 0.684 | |
| ETTh1 | 96 | 0.463 | 0.459 | 0.448 | 0.460 | 0.458 | 0.456 | 0.790 | 0.586 | 0.492 | 0.495 | 0.516 | 0.485 | 0.861 | 0.628 | 0.512 | 0.499 | 0.613 | 0.552 | 0.918 | 0.639 | 1.112 | 0.806 | 1.298 | 0.838 |
| 192 | 0.482 | 0.487 | 0.484 | 0.483 | 0.570 | 0.516 | 0.837 | 0.609 | 0.565 | 0.538 | 0.598 | 0.524 | 0.797 | 0.593 | 0.624 | 0.555 | 0.722 | 0.598 | 0.915 | 0.629 | 1.155 | 0.823 | 1.322 | 0.854 | |
| 336 | 0.603 | 0.543 | 0.589 | 0.540 | 0.608 | 0.535 | 0.780 | 0.575 | 0.721 | 0.622 | 0.657 | 0.550 | 0.941 | 0.648 | 0.691 | 0.574 | 0.750 | 0.619 | 0.939 | 0.644 | 1.179 | 0.832 | 1.347 | 0.870 | |
| 720 | 0.713 | 0.588 | 0.700 | 0.604 | 0.725 | 0.591 | 1.234 | 0.811 | 0.986 | 0.743 | 0.762 | 0.610 | 0.877 | 0.641 | 0.728 | 0.614 | 0.721 | 0.616 | 0.887 | 0.645 | 1.273 | 0.874 | 1.534 | 0.947 | |
| Avg | 0.565 | 0.524 | 0.556 | 0.522 | 0.590 | 0.525 | 0.910 | 0.860 | 0.691 | 0.600 | 0.633 | 0.542 | 0.869 | 0.628 | 0.639 | 0.561 | 0.702 | 0.596 | 0.915 | 0.639 | 1.180 | 0.834 | 1.375 | 0.877 | |
| ETTh2 | 96 | 0.300 | 0.360 | 0.275 | 0.326 | 0.331 | 0.374 | 0.404 | 0.435 | 0.357 | 0.411 | 0.353 | 0.389 | 0.378 | 0.409 | 0.382 | 0.416 | 0.413 | 0.451 | 0.389 | 0.411 | 0.678 | 0.619 | 2.022 | 1.006 |
| 192 | 0.372 | 0.371 | 0.374 | 0.373 | 0.402 | 0.411 | 0.470 | 0.474 | 0.569 | 0.519 | 0.403 | 0.414 | 0.490 | 0.467 | 0.478 | 0.474 | 0.474 | 0.477 | 0.473 | 0.455 | 0.785 | 0.666 | 2.329 | 1.104 | |
| 336 | 0.389 | 0.413 | 0.406 | 0.429 | 0.406 | 0.433 | 0.489 | 0.485 | 0.671 | 0.572 | 0.426 | 0.441 | 0.537 | 0.494 | 0.504 | 0.501 | 0.547 | 0.543 | 0.507 | 0.480 | 0.839 | 0.694 | 2.453 | 1.122 | |
| 720 | 0.403 | 0.426 | 0.427 | 0.449 | 0.449 | 0.464 | 0.593 | 0.538 | 0.824 | 0.648 | 0.477 | 0.480 | 0.510 | 0.491 | 0.499 | 0.509 | 0.516 | 0.523 | 0.477 | 0.472 | 1.273 | 0.874 | 3.816 | 1.407 | |
| Avg | 0.366 | 0.392 | 0.370 | 0.394 | 0.397 | 0.421 | 0.489 | 0.483 | 0.605 | 0.538 | 0.415 | 0.431 | 0.479 | 0.465 | 0.466 | 0.475 | 0.488 | 0.499 | 0.462 | 0.455 | 0.894 | 0.713 | 2.655 | 1.160 | |
| ETTm1 | 96 | 0.353 | 0.390 | 0.346 | 0.388 | 0.390 | 0.404 | 0.709 | 0.556 | 0.352 | 0.392 | 0.410 | 0.419 | 0.583 | 0.501 | 0.578 | 0.518 | 0.774 | 0.614 | 0.761 | 0.568 | 0.911 | 0.688 | 0.921 | 0.682 |
| 192 | 0.368 | 0.403 | 0.373 | 0.416 | 0.429 | 0.423 | 0.717 | 0.548 | 0.382 | 0.412 | 0.437 | 0.434 | 0.630 | 0.528 | 0.617 | 0.546 | 0.754 | 0.592 | 0.781 | 0.574 | 0.955 | 0.703 | 0.957 | 0.701 | |
| 336 | 0.417 | 0.428 | 0.413 | 0.426 | 0.469 | 0.439 | 0.735 | 0.575 | 0.419 | 0.434 | 0.476 | 0.454 | 0.725 | 0.568 | 0.998 | 0.775 | 0.869 | 0.677 | 0.803 | 0.587 | 0.991 | 0.719 | 0.998 | 0.716 | |
| 720 | 0.473 | 0.468 | 0.485 | 0.476 | 0.569 | 0.498 | 0.752 | 0.584 | 0.490 | 0.477 | 0.681 | 0.556 | 0.769 | 0.549 | 0.693 | 0.579 | 0.810 | 0.630 | 0.844 | 0.581 | 1.062 | 0.747 | 1.007 | 0.719 | |
| Avg | 0.402 | 0.422 | 0.404 | 0.427 | 0.464 | 0.441 | 0.728 | 0.565 | 0.411 | 0.429 | 0.501 | 0.466 | 0.677 | 0.537 | 0.722 | 0.605 | 0.802 | 0.628 | 0.797 | 0.578 | 0.980 | 0.714 | 0.971 | 0.705 | |
| ETTm2 | 96 | 0.140 | 0.242 | 0.177 | 0.261 | 0.188 | 0.269 | 0.245 | 0.322 | 0.213 | 0.303 | 0.191 | 0.274 | 0.212 | 0.285 | 0.291 | 0.399 | 0.352 | 0.454 | 0.229 | 0.308 | 0.331 | 0.430 | 0.813 | 0.688 |
| 192 | 0.207 | 0.293 | 0.241 | 0.314 | 0.251 | 0.309 | 0.274 | 0.338 | 0.278 | 0.345 | 0.252 | 0.317 | 0.270 | 0.323 | 0.307 | 0.379 | 0.694 | 0.691 | 0.291 | 0.343 | 0.400 | 0.464 | 1.008 | 0.768 | |
| 336 | 0.264 | 0.331 | 0.274 | 0.310 | 0.307 | 0.346 | 0.361 | 0.394 | 0.338 | 0.385 | 0.306 | 0.353 | 0.323 | 0.353 | 0.543 | 0.559 | 2.408 | 1.407 | 0.348 | 0.376 | 0.469 | 0.498 | 1.031 | 0.775 | |
| 720 | 0.381 | 0.387 | 0.417 | 0.390 | 0.426 | 0.417 | 0.467 | 0.442 | 0.436 | 0.440 | 0.433 | 0.427 | 0.474 | 0.449 | 0.712 | 0.614 | 1.913 | 1.166 | 0.461 | 0.438 | 0.589 | 0.557 | 1.096 | 0.791 | |
| Avg | 0.248 | 0.313 | 0.277 | 0.323 | 0.293 | 0.335 | 0.336 | 0.373 | 0.316 | 0.368 | 0.296 | 0.343 | 0.320 | 0.353 | 0.463 | 0.488 | 1.342 | 0.930 | 0.332 | 0.366 | 0.447 | 0.487 | 0.987 | 0.756 | |

### v2

@@ Appendix D Full Few-shot Forecasting Results
| Methods | IP-LLM | Time-LLM(G) | OFA | iTrans. | Dlinear | PatchTST | TimesNet | FEDformer | Autoformer | Stationary | ETSformer | LightTS | |||||||||||||
| Data. | Hori. | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE |
| Weather | 96 | 0.159 | 0.210 | 0.160 | 0.213 | 0.163 | 0.215 | 0.253 | 0.307 | 0.171 | 0.224 | 0.165 | 0.215 | 0.184 | 0.230 | 0.188 | 0.253 | 0.221 | 0.297 | 0.192 | 0.234 | 0.199 | 0.272 | 0.217 | 0.269 |
| 192 | 0.200 | 0.251 | 0.204 | 0.254 | 0.210 | 0.254 | 0.292 | 0.328 | 0.215 | 0.263 | 0.210 | 0.257 | 0.245 | 0.283 | 0.250 | 0.304 | 0.270 | 0.322 | 0.269 | 0.295 | 0.279 | 0.332 | 0.259 | 0.304 | |
| 336 | 0.257 | 0.293 | 0.255 | 0.291 | 0.256 | 0.292 | 0.322 | 0.346 | 0.258 | 0.299 | 0.259 | 0.297 | 0.305 | 0.321 | 0.312 | 0.346 | 0.320 | 0.351 | 0.370 | 0.357 | 0.356 | 0.386 | 0.303 | 0.334 | |
| 720 | 0.317 | 0.335 | 0.329 | 0.345 | 0.321 | 0.339 | 0.365 | 0.374 | 0.320 | 0.346 | 0.332 | 0.346 | 0.381 | 0.371 | 0.387 | 0.393 | 0.390 | 0.396 | 0.441 | 0.405 | 0.437 | 0.448 | 0.377 | 0.382 | |
| Avg | 0.233 | 0.272 | 0.237 | 0.275 | 0.238 | 0.275 | 0.308 | 0.338 | 0.241 | 0.283 | 0.242 | 0.279 | 0.279 | 0.301 | 0.284 | 0.324 | 0.300 | 0.342 | 0.318 | 0.323 | 0.318 | 0.360 | 0.289 | 0.322 | |
| Electricity | 96 | 0.143 | 0.243 | 0.137 | 0.240 | 0.139 | 0.237 | 0.154 | 0.257 | 0.150 | 0.253 | 0.140 | 0.238 | 0.299 | 0.373 | 0.231 | 0.323 | 0.261 | 0.348 | 0.420 | 0.466 | 0.599 | 0.587 | 0.350 | 0.425 |
| 192 | 0.159 | 0.258 | 0.159 | 0.258 | 0.156 | 0.252 | 0.171 | 0.272 | 0.164 | 0.264 | 0.160 | 0.255 | 0.305 | 0.379 | 0.261 | 0.356 | 0.338 | 0.406 | 0.411 | 0.459 | 0.620 | 0.598 | 0.376 | 0.448 | |
| 336 | 0.170 | 0.269 | 0.181 | 0.278 | 0.175 | 0.270 | 0.196 | 0.295 | 0.181 | 0.282 | 0.180 | 0.276 | 0.319 | 0.391 | 0.360 | 0.445 | 0.410 | 0.474 | 0.434 | 0.473 | 0.662 | 0.619 | 0.428 | 0.485 | |
| 720 | 0.230 | 0.315 | 0.232 | 0.317 | 0.233 | 0.317 | 0.263 | 0.348 | 0.223 | 0.321 | 0.241 | 0.323 | 0.369 | 0.426 | 0.530 | 0.585 | 0.715 | 0.685 | 0.510 | 0.521 | 0.757 | 0.664 | 0.611 | 0.597 | |
| Avg | 0.175 | 0.271 | 0.177 | 0.273 | 0.176 | 0.269 | 0.196 | 0.293 | 0.180 | 0.280 | 0.180 | 0.273 | 0.323 | 0.392 | 0.346 | 0.427 | 0.431 | 0.478 | 0.444 | 0.480 | 0.660 | 0.617 | 0.441 | 0.489 | |
| Traffic | 96 | 0.403 | 0.293 | 0.406 | 0.295 | 0.414 | 0.297 | 0.448 | 0.329 | 0.419 | 0.298 | 0.403 | 0.289 | 0.719 | 0.416 | 0.639 | 0.400 | 0.672 | 0.405 | 1.412 | 0.802 | 1.643 | 0.855 | 1.157 | 0.636 |
| 192 | 0.412 | 0.295 | 0.416 | 0.300 | 0.426 | 0.301 | 0.487 | 0.360 | 0.434 | 0.305 | 0.415 | 0.296 | 0.748 | 0.428 | 0.637 | 0.416 | 0.727 | 0.424 | 1.419 | 0.806 | 1.641 | 0.854 | 1.207 | 0.661 | |
| 336 | 0.427 | 0.316 | 0.430 | 0.309 | 0.434 | 0.303 | 0.514 | 0.372 | 0.449 | 0.313 | 0.426 | 0.304 | 0.853 | 0.471 | 0.655 | 0.427 | 0.749 | 0.454 | 1.443 | 0.815 | 1.711 | 0.878 | 1.334 | 0.713 | |
| 720 | 0.469 | 0.325 | 0.467 | 0.324 | 0.487 | 0.337 | 0.532 | 0.383 | 0.484 | 0.336 | 0.474 | 0.331 | 1.485 | 0.825 | 0.722 | 0.456 | 0.847 | 0.499 | 1.539 | 0.837 | 2.660 | 1.157 | 1.292 | 0.726 | |
| Avg | 0.427 | 0.307 | 0.429 | 0.307 | 0.440 | 0.310 | 0.495 | 0.361 | 0.447 | 0.313 | 0.430 | 0.305 | 0.951 | 0.535 | 0.663 | 0.425 | 0.749 | 0.446 | 1.453 | 0.815 | 1.914 | 0.936 | 1.248 | 0.684 | |
| ETTh1 | 96 | 0.481 | 0.474 | 0.720 | 0.533 | 0.458 | 0.456 | 0.790 | 0.586 | 0.492 | 0.495 | 0.516 | 0.485 | 0.861 | 0.628 | 0.512 | 0.499 | 0.613 | 0.552 | 0.918 | 0.639 | 1.112 | 0.806 | 1.298 | 0.838 |
| 192 | 0.518 | 0.491 | 0.747 | 0.545 | 0.570 | 0.516 | 0.837 | 0.609 | 0.565 | 0.538 | 0.598 | 0.524 | 0.797 | 0.593 | 0.624 | 0.555 | 0.722 | 0.598 | 0.915 | 0.629 | 1.155 | 0.823 | 1.322 | 0.854 | |
| 336 | 0.664 | 0.570 | 0.793 | 0.551 | 0.608 | 0.535 | 0.780 | 0.575 | 0.721 | 0.622 | 0.657 | 0.550 | 0.941 | 0.648 | 0.691 | 0.574 | 0.750 | 0.619 | 0.939 | 0.644 | 1.179 | 0.832 | 1.347 | 0.870 | |
| 720 | 0.711 | 0.584 | 0.880 | 0.584 | 0.725 | 0.591 | 1.234 | 0.811 | 0.986 | 0.743 | 0.762 | 0.610 | 0.877 | 0.641 | 0.728 | 0.614 | 0.721 | 0.616 | 0.887 | 0.645 | 1.273 | 0.874 | 1.534 | 0.947 | |
| Avg | 0.593 | 0.529 | 0.785 | 0.553 | 0.590 | 0.525 | 0.910 | 0.860 | 0.691 | 0.600 | 0.633 | 0.542 | 0.869 | 0.628 | 0.639 | 0.561 | 0.702 | 0.596 | 0.915 | 0.639 | 1.180 | 0.834 | 1.375 | 0.877 | |
| ETTh2 | 96 | 0.354 | 0.400 | 0.334 | 0.381 | 0.331 | 0.374 | 0.404 | 0.435 | 0.357 | 0.411 | 0.353 | 0.389 | 0.378 | 0.409 | 0.382 | 0.416 | 0.413 | 0.451 | 0.389 | 0.411 | 0.678 | 0.619 | 2.022 | 1.006 |
| 192 | 0.401 | 0.423 | 0.430 | 0.438 | 0.402 | 0.411 | 0.470 | 0.474 | 0.569 | 0.519 | 0.403 | 0.414 | 0.490 | 0.467 | 0.478 | 0.474 | 0.474 | 0.477 | 0.473 | 0.455 | 0.785 | 0.666 | 2.329 | 1.104 | |
| 336 | 0.442 | 0.450 | 0.449 | 0.458 | 0.406 | 0.433 | 0.489 | 0.485 | 0.671 | 0.572 | 0.426 | 0.441 | 0.537 | 0.494 | 0.504 | 0.501 | 0.547 | 0.543 | 0.477 | 0.472 | 0.839 | 0.694 | 2.453 | 1.122 | |
| 720 | 0.480 | 0.486 | 0.485 | 0.490 | 0.449 | 0.464 | 0.593 | 0.538 | 0.824 | 0.648 | 0.477 | 0.480 | 0.510 | 0.491 | 0.499 | 0.509 | 0.516 | 0.523 | 0.507 | 0.480 | 1.273 | 0.874 | 3.816 | 1.407 | |
| Avg | 0.419 | 0.439 | 0.424 | 0.441 | 0.397 | 0.421 | 0.489 | 0.483 | 0.605 | 0.538 | 0.415 | 0.431 | 0.479 | 0.465 | 0.466 | 0.475 | 0.488 | 0.499 | 0.462 | 0.455 | 0.894 | 0.713 | 2.655 | 1.160 | |
| ETTm1 | 96 | 0.388 | 0.401 | 0.412 | 0.422 | 0.390 | 0.404 | 0.709 | 0.556 | 0.352 | 0.392 | 0.410 | 0.419 | 0.583 | 0.501 | 0.578 | 0.518 | 0.774 | 0.614 | 0.761 | 0.568 | 0.911 | 0.688 | 0.921 | 0.682 |
| 192 | 0.422 | 0.421 | 0.447 | 0.438 | 0.429 | 0.423 | 0.717 | 0.548 | 0.382 | 0.412 | 0.437 | 0.434 | 0.630 | 0.528 | 0.617 | 0.546 | 0.754 | 0.592 | 0.781 | 0.574 | 0.955 | 0.703 | 0.957 | 0.701 | |
| 336 | 0.456 | 0.430 | 0.497 | 0.465 | 0.469 | 0.439 | 0.735 | 0.575 | 0.419 | 0.434 | 0.476 | 0.454 | 0.725 | 0.568 | 0.998 | 0.775 | 0.869 | 0.677 | 0.803 | 0.587 | 0.991 | 0.719 | 0.998 | 0.716 | |
| 720 | 0.554 | 0.490 | 0.594 | 0.521 | 0.569 | 0.498 | 0.752 | 0.584 | 0.490 | 0.477 | 0.681 | 0.556 | 0.769 | 0.549 | 0.693 | 0.579 | 0.810 | 0.630 | 0.844 | 0.581 | 1.062 | 0.747 | 1.007 | 0.719 | |
| Avg | 0.455 | 0.435 | 0.487 | 0.461 | 0.464 | 0.441 | 0.728 | 0.565 | 0.411 | 0.429 | 0.501 | 0.466 | 0.677 | 0.537 | 0.722 | 0.605 | 0.802 | 0.628 | 0.797 | 0.578 | 0.980 | 0.714 | 0.971 | 0.705 | |
| ETTm2 | 96 | 0.192 | 0.274 | 0.224 | 0.296 | 0.188 | 0.269 | 0.245 | 0.322 | 0.213 | 0.303 | 0.191 | 0.274 | 0.212 | 0.285 | 0.291 | 0.399 | 0.352 | 0.454 | 0.229 | 0.308 | 0.331 | 0.430 | 0.813 | 0.688 |
| 192 | 0.246 | 0.313 | 0.260 | 0.317 | 0.251 | 0.309 | 0.274 | 0.338 | 0.278 | 0.345 | 0.252 | 0.317 | 0.270 | 0.323 | 0.307 | 0.379 | 0.694 | 0.691 | 0.291 | 0.343 | 0.400 | 0.464 | 1.008 | 0.768 | |
| 336 | 0.301 | 0.340 | 0.312 | 0.349 | 0.307 | 0.346 | 0.361 | 0.394 | 0.338 | 0.385 | 0.306 | 0.353 | 0.323 | 0.353 | 0.543 | 0.559 | 2.408 | 1.407 | 0.348 | 0.376 | 0.469 | 0.498 | 1.031 | 0.775 | |
| 720 | 0.400 | 0.403 | 0.424 | 0.416 | 0.426 | 0.417 | 0.467 | 0.442 | 0.436 | 0.440 | 0.433 | 0.427 | 0.474 | 0.449 | 0.712 | 0.614 | 1.913 | 1.166 | 0.461 | 0.438 | 0.589 | 0.557 | 1.096 | 0.791 | |
| Avg | 0.284 | 0.332 | 0.305 | 0.344 | 0.293 | 0.335 | 0.336 | 0.373 | 0.316 | 0.368 | 0.296 | 0.343 | 0.320 | 0.353 | 0.463 | 0.488 | 1.342 | 0.930 | 0.332 | 0.366 | 0.447 | 0.487 | 0.987 | 0.756 | |